Operations Report – November 24th

Crew 183 Operations Report 24-11-2017

SOL: 2

Name of person filing report: Ashton Stoop

Non-nominal systems:

Power system – SOC

Notes on non-nominal systems:

The power system calculates SOC based on an algorithm, it is not a measured value. This has been causing strange fluctuations on when reading the SOC.

READING FROM 24 NOV 2017

TIME     SOC READINGS

6:32       61%

9:25       69%

10:11     100%

11:40     100%

16:10     88%

18:37     100%

Generator (hours run): 14

Solar— 100%

Diesel –  62%

Propane –  42% volume

Ethanol Free Gasoline (5 Gallon containers for ATV) –   0 gallons remaining

Water (trailer) – 100 gallons added to static

Water (static) –   325 gallons

Trailer to Static Pump used –  yes

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  45

Toilet tank emptied: yes

ATV’s Used: Honda, 300, 350.1, 350.2

Oil Added? No

ATV Fuel Used:  5 Gals

# Hours the ATVs were Used today: 3

Notes on ATVs: ATV 350.1 had a bit of trouble starting up upon return from EVA. Once it got going everything seemed to be fine. We’ll keep an eye on this for the next few SOLs.

Deimos rover used:   Yes

Hours: 1

Beginning charge: 100%

Ending charge: 90%

Currently charging: Yes

 

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

 

Spirit rover used:  NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

 

Opportunity rover used:  NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? NO

General notes and comments:

The SOC readings seem off. Looking to further investigate this matter.

Summary of internet:

We hit our internet limit around 20:00.

Summary of suits and radios:

Not used

Summary of Hab operations:

Everything is in good shape. We had some bomb crab biscuits for dinner.

Summary of GreenHab operations:

A new lighting timer has been delivered and installed in the GreenHab.

Summary of ScienceDome operations:

New lights have been delivered. Tomorrow, we need to install lightbulbs and test.

Summary of RAMM operations:

Not used. It is not yet fully operational.

Summary of health and safety issues:

Four carbon monoxide detectors have been ordered.

Questions, concerns and requests to Mission Support:

-Dear mission support, how exactly does the SOC reading on the power system operate? We have had fluctuations in the SOC readings which do not correlate closely with our estimated values of power generation and power consumption (Please refer to SOC Readings table above).

-We have no additional supply of gasoline for the ATVs. All that’s left is what’s in the vehicles. We need fuel ASAP!